I not sure if I should move the y to be with the 3 in the problem: 3=Ln(x/y) if I have to solve for y.
Answer by stanbon(57377)

I not sure if I should move the y to be with the 3 in the problem: 3=Ln(x/y) if I have to solve for y.
-------------------------------------
3=Ln(x/y)
----
x/y = e^3
---
y = x/e^3
=================
Cheers,
Stan H.